I am not getting the right answer can you show me the process to obtain the answer?
y+2=5(x+1) for y

Found 2 solutions by josgarithmetic, mananth:
Answer by josgarithmetic(39629)

(1) Distributive Property on the right member
(2) Additive Inverse of +2, apply to BOTH members
(3) Combine any like-terms

Answer by mananth(16946)

y+2=5(x+1)
y+2-2=5(x+1)-2
y=5(x+1)-2
y=5x+5-2

y=5x+3